OpenStudy (anonymous):

The sum of the squares of 2 consecutive negative integers is 41. What are the numbers? Which of the following equations is the result of using the factoring method to solve the problem? (n - 5)(n - 4) = 0 (n - 5)(n + 4) = 0 (n + 5)(n - 4) = 0 (n + 5)(n + 4) = 0

OpenStudy (anonymous):

ok maybe try \[n^2+(n+1)^2=41\] and see what happens

OpenStudy (anonymous):

I got (n+5)(n-4)

OpenStudy (anonymous):

multiplying out gives \[n^2+n^2+2n+1=41\] subtract and get \[2n^2+2n-40=0\] divide by 2 and get \[n^2+n-20=0\] then \[(n+5)(n-4)=0\]

OpenStudy (anonymous):

yeah i got the same as you